OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 128 yd^2. The length is __ yards. The width is __ yards.

OpenStudy (anonymous):

2L x w = 128 should be your model equation

OpenStudy (anonymous):

x = width x*2x = 2x^2 = 128 x = 8 The length is 16 yards The width is 8 Yards

OpenStudy (anonymous):

he's quick

OpenStudy (anonymous):

agreed lol

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

is the 128^2 yards though?

OpenStudy (anonymous):

those are just your units

OpenStudy (anonymous):

Yes, 16*8 = 128, I didn't include units.

OpenStudy (anonymous):

sweet! Thanks